A psychologist studying addiction tests whether cravings for cocaine and relapse are independent. The following table lists the observed frequencies in the small sample of people who use drugs. Obs. Freq. Relapse Yes No Cravings Yes 22 10 32 No 9 14 23 31 24 N = 55 (a) Conduct a chi-square test for independence at a 0.05 level of significance. (Round your answer to two decimal places.) $ \chi_{obt}^2 $ = Incorrect: Your answer is incorrect. Decide whether to retain or reject the null hypothesis. Retain the null hypothesis. Reject the null hypothesis. (b) Compute effect size using ϕ and Cramer's V. Hint: Both should give the same estimate of effect size. (If necessary, round your intermediate steps to two or more decimal places. Round your answers to two decimal places.) ϕ = V =
Answer #1

a) applying chi square test:

Expected Ei=row total*column total/grand total Yes No Total
yes 18.04 13.96 32
No 12.96 10.04 23
total 31 24 55
chi square    χ2 =(Oi-Ei)2/Ei Yes No Total
yes 0.8710 1.1251 1.996
No 1.2119 1.5653 2.777
total 2.083 2.690 4.7734

chi-square X2obt =4.77

Reject the null hypothesis

b)

Cramer's V =sqrt(X2/n) =0.29